OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 a.(-5, 2) b.(5, -2) c.(-5, -2) d.(5, 2)

OpenStudy (mertsj):

Why don't you try replacing x and y with the values of the ordered pairs and see which makes a true statement?

OpenStudy (anonymous):

Just see for second inequality... by putting values and check that it must not imply the first equality y=x-5

OpenStudy (mertsj):

\[y \le x-5\] \[2\le -5-2 \] \[2\le-7\] False Move on to the next point

OpenStudy (mertsj):

\[y \le x-5\] \[-2\le 5-2\] \[-2\le 3\]

OpenStudy (mertsj):

True. so try that same point in the second inequality

OpenStudy (mertsj):

\[y \ge -x-4\] \[-2\ge -5-4\] \[-2\ge -9\] True. Since that ordered pair makes both sentences true, that must be the answer.